A coin bank contains twenty-four nickels and quarters. The number of quarters in the bank is six less than twice the number of nickels. Find the number of nickels and quarters in the bank. How do I do this question?
Answer by mananth(16949)

A coin bank contains twenty-four nickels and quarters. The number of quarters in the bank is six less than twice the number of nickels. Find the number of nickels and quarters in the bank. How do I do this question?
...
nickels = n
quarters = 2n-6
...
Total = 24 coins
n+2n-6 =24
3n-6 =24
3n=30
n=10
quarters = 2n-6 = 2*10-6=14
Quarters = 14 numbers
..
m.ananth@hotmail.ca